OpenStudy (anonymous):

A circle has a central angle of 40 degrees. The radius of the circle is 3 meters. Find the area of the sector created by the central angle. Leave pi in your answer.

OpenStudy (anonymous):

1/2 r^2 x angle = 1/2 x (3 metres)^2 x 40 degrees = 1/2 x 9m^2 x 40 degrees = 4.5m^2 x 40 degrees = 180 m^2

OpenStudy (anonymous):

okay